OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2. simplify (1/2A + 1/2B) and What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (anonymous):

someone please help me with this one.

OpenStudy (anonymous):

A= 3/6 = 1/2 B = 3/6 = 1/2 So, \[(1/2)^2 + (1/2)^2 = 2/4 = 1/2\] = 1/2

OpenStudy (anonymous):

how do I get the answer to B?

OpenStudy (anonymous):

Options are??

OpenStudy (anonymous):

What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?